OpenStudy (anonymous):

At a hospital emergency room, a doctor has determined that for every 10 patients injured from an automobile wreck, 7 were not wearing their seat belts. If the emergency room saw 500 patients last month that were injured from an automobile wreck, what is the best prediction for how many of them were not wearing their seat belt?

OpenStudy (anonymous):

7/10=x/500 7x500=3500 3500/10=350